# Multipartite Embezzlement of Entanglement

### v1

@@ 2 Overview of results and discussion
We emphasize that since we consider embezzlement relative to local operations and not relative to local operations and classical communication (LOCC), the above Lemma does not follow from the fact that any pure finite dimensional entangled state can be obtained from a sufficiently high dimensional maximally entangled state by LOCC. In particular, it is, in fact, not sufficient if the embezzling family allows to embezzle Bell states (see Remark 31).
For bipartite entanglement, [long_paper] establishes precise quantitative criteria on an operator algebraic level for the existence of embezzling states on bipartite quantum systems. In particular, type ${\mathrm{III}}_{1}$ von Neumann algebras are singled out as the local von Neumann algebras of universal embezzlers.
@@ 3 Multipartite embezzling families and states
If a bipartite system $({\mathcal{M}}_{1},{\mathcal{M}}_{2})$ on a Hilbert space ${\mathcal{H}}$ is split then we can find von Neumann algebras ${\mathcal{N}}_{j}\cong{\mathcal{M}}_{j}$ on Hilbert spaces ${\mathcal{H}}_{j}$, $j=1,2$, and a unitary isomorphism ${\mathcal{H}}\cong{\mathcal{H}}_{1}\otimes{\mathcal{H}}_{2}$ such that
$({\mathcal{M}}_{1},{\mathcal{M}}_{2})\cong({\mathcal{N}}_{1}\otimes 1,1\otimes{\mathcal{N}}_{2}).$ (10)
@@
Let $({\mathcal{M}}_{1},{\mathcal{M}}_{2},{\mathcal{M}}_{3})$ be a tripartite system on ${\mathcal{H}}$ fulfilling Haag duality. If the bipartite system $({\mathcal{M}}_{1},{\mathcal{M}}_{3})$ is split, then there exists a unitary $u:{\mathcal{H}}\to{\mathcal{H}}_{1}\otimes{\mathcal{H}}_{3}$ inducing an isomorphism
$({\mathcal{M}}_{1},{\mathcal{M}}_{2},{\mathcal{M}}_{3})\cong({\mathcal{N}}_{1}\otimes 1,{\mathcal{N}}_{1}^{\prime}\otimes{\mathcal{N}}_{3}^{\prime},1\otimes{\mathcal{N}}_{3}),$ (11)
where ${\mathcal{N}}_{1}$ and ${\mathcal{N}}_{2}$ are the von Neumann algebras defined by ${\mathcal{N}}_{1}\otimes 1=u{\mathcal{M}}_{1}u^{*}$ and $1\otimes{\mathcal{N}}_{3}=u{\mathcal{M}}_{3}u^{*}$.
@@
We apply (10), and use Haag duality to find ${\mathcal{M}}_{2}=({\mathcal{M}}_{1}\vee{\mathcal{M}}_{3})^{\prime}\cong({\mathcal{N}}_{1}\otimes{\mathcal{N}}_{3})^{\prime}={\mathcal{N}}_{1}^{\prime}\otimes{\mathcal{N}}_{3}^{\prime}$. ∎
@@ 3.4 The limit of the Leung-Toner-Watrous embezzling family
The Hilbert space ${\mathcal{H}}$ carries an $N$-partite structure in the sense that the observables of each party $x\in\{1,\ldots,N\}$ are described by a factor ${\mathcal{M}}_{x}$ acting on ${\mathcal{H}}$.  [^11] The factor ${\mathcal{M}}_{x}$ is defined as the weak closure of $\bigcup_{n\in{\mathbb{N}}}\left(\bigotimes_{k=1}^{n}1_{[1,x-1]}\otimes M_{D_{k}}({\mathbb{C}})\otimes 1_{[x+1,N]}\right)$ in its natural action on ${\mathcal{H}}$. By construction, ${\mathcal{M}}_{x}$ is isomorphic with the ITPFI factor
@@ 4 Monopartite and bipartite embezzling families
To obtain our results, we first show that it suffices for a monopartite embezzling family to be able to embezzle arbitrary maximally mixed states in order to embezzle arbitrary mixed states. There are two ways one can show this: First, the results in [long_paper] based on the flow of weights imply that for (monopartite) embezzling states (instead of families) being able to embezzle maximally mixed states of dimension $d_{1}$ and $d_{2}$ such that $\log(d_{1})/\log(d_{2})$ is irrational is sufficient to be able to embezzle arbitrary mixed states. By the monopartite analog of Theorem 20, having a consistent monopartite family that can embezzle arbitrary maximally mixed states yields a state that can embezzle arbitrary maximally mixed states. Hence, it is a monopartite embezzling state, and its restriction (which yields again the starting family) yields a monopartite embezzling family.
@@ 4.2 A simpler Leung-Toner-Watrous embezzling family
In section 3.4, we showed that one can take the limit of the LTW construction to obtain a multipartite, universal embezzler. One obvious disadvantage of the construction is its ‘inefficiency’: Tensoring $\varepsilon$-covers of high-dimensional unit balls seems too wasteful in terms of Hilbert space dimension. Here, we consider a simpler version of the LTW family in the monopartite setting and show that it yields a universal monopartite embezzler. Via purification, the construction naturally yields a universal embezzler for bipartite entanglement.
Instead of relying on abstract arguments, we will directly compute the asymptotic ratio set $r_{\infty}({\mathcal{M}})$ of the resulting von Neumann algebra, which coincides with the Connes invariant $S({\mathcal{M}})$. By showing that $r_{\infty}({\mathcal{M}})={\mathbb{R}}_{+}$, it follows that ${\mathcal{M}}$ has type III1 [araki1968factors].
@@
with $u\in M_{d}({\mathbb{C}})^{\otimes n}$ being the unitary that cyclically permutes the tensor factors [^13] Note that we use approximate catalysts of a slightly different form compared to Section 3.4..
@@ Appendix A Split bipartite systems admit no embezzling states
In this appendix we prove a no-go theorem for bipartite embezzlement in split bipartite systems. In fact we will rule out the possibility to embezzle a single pure entangled state.
@@
As a biproduct of our proof, we obtain that any system which admits embezzlement of a single entangled state must be type III.

### v2

@@ Acknowledgments.
Declarations.
The authors have no conflicts of interest to declare. No data was generated or processed in this work.
@@ 2 Overview of results and discussion
We emphasize that since we consider embezzlement relative to local operations and not relative to local operations and classical communication (LOCC), the above Lemma does not follow from the fact that any pure finite-dimensional entangled state can be obtained from a sufficiently high-dimensional maximally entangled state by LOCC. In particular, it is, in fact, not sufficient if the embezzling family allows to embezzle Bell states (see Remark 31).
For bipartite entanglement, [long_paper] establishes precise quantitative criteria on an operator algebraic level for the existence of embezzling states on bipartite quantum systems. In particular, type ${\mathrm{III}}_{1}$ von Neumann algebras are singled out as the local von Neumann algebras of universal embezzlers. In contrast, in bipartite systems composed of commuting factors in Haag duality, all pure states are embezzling states relative to LOCC if and only if the factors have type ${\mathrm{III}}$, irrespective of the subtype [van_luijk_pure_2024].
@@ 3 Multipartite embezzling families and states
If a bipartite system $({\mathcal{M}}_{1},{\mathcal{M}}_{2})$ on a Hilbert space ${\mathcal{H}}$ is split then we can find a tensor product decomposition ${\mathcal{H}}={\mathcal{H}}_{1}\otimes{\mathcal{H}}_{2}$ and von Neumann algebras ${\mathcal{N}}_{j}\cong{\mathcal{M}}_{j}$ on ${\mathcal{H}}_{j}$, $j=1,2$, such that [^3] More precisely, there is a unitary $u:{\mathcal{H}}\to{\mathcal{H}}_{1}\otimes{\mathcal{H}}_{2}$ such that $u^{*}{\mathcal{M}}_{1}u={\mathcal{N}}_{1}\otimes 1$ and $u^{*}{\mathcal{M}}_{2}u=1\otimes{\mathcal{N}}_{2}$.
$({\mathcal{M}}_{1},{\mathcal{M}}_{2})=({\mathcal{N}}_{1}\otimes 1,1\otimes{\mathcal{N}}_{2}).$ (10)
@@
Let $({\mathcal{M}}_{1},{\mathcal{M}}_{2},{\mathcal{M}}_{3})$ be a tripartite system on ${\mathcal{H}}$ fulfilling Haag duality. If the bipartite system $({\mathcal{M}}_{1},{\mathcal{M}}_{3})$ is split, then there exists a tensor product decomposition ${\mathcal{H}}={\mathcal{H}}_{L}\otimes{\mathcal{H}}_{R}$ and von Neumann algebras ${\mathcal{N}}_{1}\cong{\mathcal{M}}_{1}$ on ${\mathcal{H}}_{L}$, ${\mathcal{N}}_{2}\cong{\mathcal{M}}_{2}$ on ${\mathcal{H}}_{R}$ such that
[…]
@@
We apply (10), and use Haag duality to find ${\mathcal{M}}_{2}=({\mathcal{M}}_{1}\vee{\mathcal{M}}_{3})^{\prime}=({\mathcal{N}}_{1}\otimes{\mathcal{N}}_{3})^{\prime}={\mathcal{N}}_{1}^{\prime}\otimes{\mathcal{N}}_{3}^{\prime}$. ∎
@@ 3.4 The limit of the Leung-Toner-Watrous embezzling family
The Hilbert space ${\mathcal{H}}$ carries an $N$-partite structure in the sense that the observables of each party $x\in\{1,\ldots,N\}$ are described by a factor ${\mathcal{M}}_{x}$ acting on ${\mathcal{H}}$ [^12] The factor ${\mathcal{M}}_{x}$ is defined as the weak closure of $\bigcup_{n\in{\mathbb{N}}}\left(\bigotimes_{k=1}^{n}1_{[1,x-1]}\otimes M_{D_{k}}({\mathbb{C}})\otimes 1_{[x+1,N]}\right)$ in its natural action on ${\mathcal{H}}$. By construction, ${\mathcal{M}}_{x}$ is isomorphic with the ITPFI factor
@@ 4 Monopartite and bipartite embezzling families
To obtain our results, we first show that it suffices for a monopartite embezzling family to be able to embezzle arbitrary maximally mixed states in order to embezzle arbitrary mixed states. There are two ways one can show this: First, the results in [long_paper] based on the flow of weights imply that for (monopartite) embezzling states (instead of families) being able to embezzle maximally mixed states of dimension $d_{1}$ and $d_{2}$ such that $\log(d_{1})/\log(d_{2})$ is irrational is sufficient to be able to embezzle arbitrary mixed states. Second, by the monopartite analog of Theorem 20, having a consistent monopartite family that can embezzle arbitrary maximally mixed states yields a state that can embezzle arbitrary maximally mixed states. Hence, it is a monopartite embezzling state, and its restriction (which yields again the starting family) yields a monopartite embezzling family.
@@ 4.2 A simpler Leung-Toner-Watrous embezzling family
In section 3.4, we showed that one can take the limit of the LTW construction to obtain a multipartite, universal embezzler. One obvious disadvantage of the construction is its ‘inefficiency’: Tensoring $\varepsilon$-covers of high-dimensional unit balls seems too wasteful in terms of Hilbert space dimension. Here, we consider a simpler version of the LTW family in the monopartite setting and show that it yields a universal monopartite embezzler. Via purification, the construction naturally yields a universal embezzler for bipartite entanglement. One way to quantify the efficiency of an embezzling family of states on finite spin chains is in terms of the asymptotic scaling of the circuit complexity. This is considered in [schwartzman_complexity_2024], where the author proves a lower bound on the complexity of all embezzling families.
In this subsection, we do two things: We note a simpler version of the LTW family, which is defined in the same way as the LTW family but only takes maximally entangled states. Moreover, we give direct proof of the claim that the limiting system is of type ${\mathrm{III}}_{1}$ by directly computing the asymptotic ratio set $r_{\infty}({\mathcal{M}})$ of the resulting von Neumann algebra.
@@
with $u\in M_{d}({\mathbb{C}})^{\otimes n}$ being the unitary that cyclically permutes the tensor factors. [^14] Note that we use approximate catalysts of a slightly different form compared to Section 3.4.
@@ Appendix A Split bipartite systems admit no embezzling states
In this appendix, we prove a no-go theorem for bipartite embezzlement in split bipartite systems. In fact, we will rule out the possibility to embezzle a single pure entangled state.
@@
As a byproduct of our proof, we obtain that any system which admits embezzlement of a single entangled state must be type III.